# Autoregressive Language Model Pre-training

### v1

@@ Abstract
Mixture-of-experts (MoE) models facilitate efficient scaling; however, training the router network introduces the challenge of optimizing a non-differentiable, discrete objective. Recently, a fully-differentiable MoE architecture, SMEAR, was proposed (Muqeeth et al., 2023), which softly merges experts in the parameter space; nevertheless, its effectiveness was only demonstrated in downstream fine-tuning on classification tasks. In this paper, we present Lory, the first approach that scales such architectures to autoregressive language model pre-training. Lory introduces two key techniques: (1) a causal segment routing strategy that achieves high efficiency for expert merging operations while preserving the autoregressive nature of language models; (2) a similarity-based data batching method that encourages expert specialization by grouping similar documents in training instances. We pre-train a series of Lory models on 150B tokens from scratch, with up to 32 experts and 30B (1.5B active) parameters. Experimental results show significant performance gains over parameter-matched dense models on both perplexity (+13.9%) and a variety of downstream tasks (+1.5%-11.1%). Despite segment-level routing, Lory models achieve competitive performance compared to state-of-the-art MoE models with token-level routing. We further demonstrate that the trained experts in Lory capture domain-level specialization without supervision. Our work highlights the potential of fully-differentiable MoE architectures for language model pre-training and advocates future research in this area.
@@ 1 Introduction
Mixture-of-experts (MoE) architectures with sparse activation enable the scaling of model sizes while maintaining high training and inference efficiency (Lepikhin et al., 2021; Fedus et al., 2022; Du et al., 2022; Zoph et al., 2022; Lewis et al., 2021; Zhou et al., 2022; Jiang et al., 2024; Xue et al., 2024; Shen et al., 2024). However, training the routing network in MoE architectures introduces the challenge of optimizing a non-differentiable, discrete objective (Shazeer et al., 2017; Zoph et al., 2022). Various techniques—such as switch routing (Fedus et al., 2022), top-$k$ expert-choice routing (Zhou et al., 2022), and linear programming (Lewis et al., 2021)—have been developed to address this challenge, often requiring carefully designed load balancing objectives (Fedus et al., 2022) or introducing additional complexity in assignment algorithms (Lewis et al., 2021; Roller et al., 2021).
@@
Recent research has started to explore fully-differentiable MoE architectures as an alternative to overcome training difficulty. Notably, SMEAR (Muqeeth et al., 2023) is an approach that softly merges experts as a weighted average of all the experts’ parameters, as opposed to activating the top-$k$ experts. However, the effectiveness of SMEAR has only been demonstrated in small-scale fine-tuning experiments on downstream classification tasks (Wang et al., 2018). In this work, we propose Lory [^1] Lory is a tribe of parrots with rainbow-like colors, which resembles the spirit of ‘soft’ MoE., the first approach that scales such fully-differentiated MoE architectures to autoregressive language model pre-training. Unlike text classification tasks which only require routing each input sequence to different experts, language modeling makes predictions for each input token, and performing token-level routing is prohibitively expensive as the computational cost of merging operations scales linearly with the number of experts.
Lory is based on two key techniques (Figure 1). We first propose causal segment routing. For a sequence of input tokens, we split them into multiple segments with a fixed length, and use the previous segment to determine the router’s weights and calculate the merged expert for the subsequent segment. During inference, we can simply use the prompt to make a single routing decision throughout the generation. This segment-level routing strategy preserves the autoregressive nature of language models, while keeping the merging operations efficient. However, since the text data for pre-training language models usually concatenates random sets of documents, we find that such routing can lead to scenarios in which experts are not sufficiently specialized. Hence, we propose our second technique—similarity-based data batching for MoE training, which groups semantically similar documents to form consecutive segments. This idea has been recently proposed to train LMs to better reason across document boundaries (Shi et al., 2024), while we find that it leads to more effective training of expert routing.
We pre-train a series of Lory models from scratch under a training budget of 150B tokens, with 0.3B and 1.5B active parameters, and 8, 16 or 32 experts (up to 6.8B and 29.5B full parameters; see Table 3). Experimental results show that our Lory models significantly outperform equal-sized dense models trained with the same amount of data, achieving performance gains on both perplexity ($+13.9\%$), and a wide range of downstream tasks including commonsense reasoning ($+3.7\%$), reading comprehension ($+3.3\%$), closed-book QA ($+1.5\%$), and text classification ($+11.1\%$). Interestingly, despite that Lory uses segment-level routing, we find it achieves competitive performance compared to state-of-the-art MoE models with token-level, non-differentiable discrete routing (Zhou et al., 2022). Our analysis further shows that the trained experts in Lory capture domain-level specialization without any supervision, making it distinct from previous MoE LMs with token-level routing, which only exhibits local patterns uniformly distributed across different domains (Xue et al., 2024; Jiang et al., 2024). Together, we present the first fully-differentiated MoE model that is suitable for language model pre-training, and demonstrate its effectiveness at scale. We hope our work sheds light on the potential of fully differentiable MoE architectures in cultivating specialized experts and we seek to encourage continued exploration in this research field.
@@ 2.1 Sparsely-activated MoE
Transformer-based MoE language models typically substitute feed-forward network (FFN) layers with sparsely-activiated MoE layers (Shazeer et al., 2017; Fedus et al., 2022; Zoph et al., 2022). Assume an MoE layer consists of $E$ expert FFNs, each parameterized as $\text{FFN}(\cdot;\theta_{1}),\dots,\text{FFN}(\cdot;\theta_{E})$, where the function $\text{FFN}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ defines a single expert module. For each token $x$ in a sequence, an MoE layer takes the hidden representation $h_{x}\in\mathbb{R}^{d}$ as the input and computes its output $o_{x}\in\mathbb{R}^{d}$ by sparsely activating $k$ experts in this layer and aggregating the outputs through a weighted sum:
@@
The routing weight $e_{i}$ for the $i$-th expert is measured by a routing network or router $R$, which takes $h_{x}$ as input and calculates the weight for each expert. In practice, to achieve sparsity and computational efficiency, only one (Fedus et al., 2022) or top-$k$ (Lepikhin et al., 2021) experts with the highest routing weights are activated at each MoE layer. The weights of the remaining experts are set to 0 (i.e., $e_{i}=0$), eliminating the need to compute $\text{FFN}(h_{x};\theta_{i})$ and effectively deactivating the $i$-th expert.
@@ 2.2 Fully Differentiable MoE Architectures via Expert Merging
The primary challenges in training sparsely activated MoE models arise from the difficulty in training discrete routers. A promising direction is to design fully differentiable MoE architectures that do not depend on extra loss formulations for stablized training. A recent model architecture (Muqeeth et al., 2023) demonstrates the feasibility by computing a weighted average of all expert FFNs in the parameter space (Matena & Raffel, 2022; Wortsman et al., 2022), thereby creating a "merged FFN". Given an input $x$ and its corresponding routing weights $e_{i}$, the output $o_{x}$ of a merged FFN is computed as:
@@
However, naively extending it to autoregressive language models, which would require computing the merged FFN for each token in a sequence, would be infeasible as the computational costs of merging operations scales linearly with the number of experts. SMEAR (Muqeeth et al., 2023) has only been evaluated for downstream fine-tuning on text classification tasks, which makes routing decisions based on a pooling representation of the entire input sequence, i.e., $e_{i}=\text{Softmax}(R(\frac{\sum_{j=1}^{L}h_{x_{j}}}{L}))_{i}$. Such operations will disrupt the autoregressive property in language model pre-training. In this work, we address these challenges by developing a fully differentiable MoE architecture suitable for autoregressive language modeling, and pre-train such models at scale.
@@
In this section, we present Lory, an approach for pre-training fully differentiable MoE language models (Figure 1). The core technique that enables Lory to be fully differentiable is expert merging (Muqeeth et al., 2023, see details in Section 2.2). To make it computationally feasible, we propose a causal segment routing method that only merges experts once for each segment, effectively reducing the number of merging operations (Section 3.1). We also propose a data batching strategy of grouping semantically similar texts, which is crucial for effective training of the segment-level router (Section 3.2).
@@
We denote an input sequence of $L$ tokens as $X=(x_{1},x_{2},\ldots,x_{L})$. By considering a segment size $T$, we divide the input sequence into $N=\lceil L/T\rceil$ segments, denoted as $S_{1},S_{2},\ldots,S_{N}$. We use $R$ to denote the routing network (parameterized as a linear layer) that computes the weights for expert merging. Let $h_{x}$ represent the hidden representation of the token $x$. The parameters of the $i$-th expert FFN are denoted by $\theta_{i}$.
@@ 3.1 Efficient Expert Merging via Causal Segment Routing
An intuitive way of reducing the computational cost is to use segment-level routing instead of token-level routing, which can reduce the number of merging operations from $L$ to $N$ times. However, simply using the current segment to compute the routing weights can cause information leakage.
@@
We propose causal segment routing to effectively route information across segments in an autoregressive manner. [^2] A piece of pseudocode of the causal segment routing strategy can be found in Appendix A. It merges FFNs in an MoE layer based on the previous segment’s information, and uses it to process the current segment. Specifically, given a training instance $X$ that consists of $L$ tokens (e.g., $L=4096$), we split the training instance into $N$ segments, each of which contains $T$ (e.g., $T=256$) consecutive tokens. For the $k$-th segment $S_{k}$ when $k>1$, we compute the average of the hidden representations of its preceding segment $S_{k-1}$ , denoted as $\bar{h}_{k-1}$. Using the average hidden representation allows the model to adapt to prompts of varying lengths during inference. $\bar{h}_{k-1}$ is then utilized to determine the routing weights, resulting in a merged expert $\bar{\theta}$:
@@
We then use the merged expert $\bar{\theta}$ to process all the tokens in the current segment $S_{k}$, i.e., $o_{x}=\text{FFN}(h_{x};\bar{\theta}),\forall x\in S_{k}$. This approach guarantees that the routing decisions made by the model are based exclusively on data from preceding positions. For the first segment $S_{1}$, the representation of the segment itself is used to compute the merging weights for its own FFN. To prevent information leakage, we implement a stop-gradient operation on $R(\bar{h}_{1})$. As demonstrated in Appendix B, merging experts at the segment level incurs minimal overhead compared to the training of dense models.
@@
During inference, we begin with a given prompt and make a single routing decision per layer based on the average hidden representations of the prompt. This routing decision determines a merged FFN and it is used consistently throughout the entire generation process. It is important to note that this inference process is as simple and computationally efficient as dense models. [^3] In Appendix G.3, we compare the prompt-only routing strategy to using the causal segment routing strategy that faithfully follows the training design. We find that these two strategies do not lead to significant differences in performance on downstream tasks. Given the efficiency advantage of making a single routing decision, we adopt the prompt-only strategy as the default approach. In Appendix H.2, we discuss the potential of converting Lory to sparsely activated MoE models for memory-efficient inference, which we leave it as future work.
@@ 3.2 Similarity-based Data Batching
The standard practice of pre-training LMs is to randomly concatenate documents to construct training instances with a fixed length. This could lead to under-specialized experts, because tokens within adjacent segments may come from very different and irrelevant documents. To mitigate this issue, we employ a similarity-based data batching technique inspired by Shi et al. (2024), which sequentially concatenates similar documents to construct training instances. This encourages high similarity between adjacent segments, enabling the experts to specialize in different domains or topics. We measure document similarity using Contriever (Izacard et al., 2022) and concatenate similar documents based on a greedy search algorithm (see Appendix C). Although we employ a data batching technique similar to Shi et al. (2024), our motivation differs from theirs. While their work aims to improve language models’ reasoning across document boundaries, we find this technique effective in encouraging expert specialization in training MoE models.
@@ 4.1 Setups
We employ the AdamW optimizer (Loshchilov & Hutter, 2019) with $\beta_{1}=0.9$ and $\beta_{2}=0.95$ and use a learning rate of 2e-4 with a cosine learning rate scheduler. All models with a batch size of 1 million tokens. We employ the data parallelism with the ZeRO optimization (Rajbhandari et al., 2020) for distributed training. [^6] In Appendix H.1, we discuss parallelism strategies when scaling up model sizes (e.g., $>100$B). At the beginning of training, we train a parameter-matched dense model and duplicate the FFN layers as initialization of the MoE model. In our experiments, we use the first 5% training steps as the warmup to initialize the MoE weights. We find that without warmup training, there may be more experts under-utilized (see Appendix G.4 for an ablation study). We also apply a linear warmup to the learning rate scheduler for the first $5\%$ training steps. We train our models with up to 64 A100 GPUs.
@@
We randomly sample a subset of the Commoncrawl dataset (Wenzek et al., 2019) as the training data. The full training dataset consists of 150 billion tokens in total. We apply the similarity-based data batching method on this subset of construct all the training instances, following Shi et al. (2024). See Appendix C for details of the data batching method.

### v2

@@ Abstract
Mixture-of-experts (MoE) models facilitate efficient scaling; however, training the routing network introduces the challenge of optimizing a non-differentiable, discrete objective. Recently, a fully differentiable MoE architecture called SMEAR was proposed (Muqeeth et al., 2023), which softly merges experts in the parameter space. Nevertheless, its effectiveness has only been demonstrated in downstream fine-tuning on classification tasks. In this work, we present Lory, a novel approach that scales such architectures to autoregressive language model pre-training. Lory introduces two key techniques: (1) a causal segment routing strategy that achieves high efficiency for expert merging operations while preserving the autoregressive nature of language models, and (2) a similarity-based data batching method that encourages expert specialization by grouping similar documents in training instances. We pre-train a series of Lory models from scratch on 150B tokens, with up to 32 experts and 30B (1.5B active) parameters. Experimental results show significant performance gains over parameter-matched dense models in both perplexity (+13.9%) and a variety of downstream tasks (+1.5%-11.1%). Despite segment-level routing, Lory models achieve competitive performance compared to state-of-the-art MoE models with token-level routing. We further demonstrate that the trained experts capture domain-level specialization without supervision. Our work highlights the potential of fully differentiable MoE architectures for language model pre-training and advocates for future research in this area.
@@ 1 Introduction
Mixture-of-experts (MoE) architectures with sparse activation enable the scaling of model sizes while maintaining high training and inference efficiency (Lepikhin et al., 2021; Fedus et al., 2022; Du et al., 2022; Zoph et al., 2022; Lewis et al., 2021; Zhou et al., 2022; Jiang et al., 2024; Xue et al., 2024; Shen et al., 2024). However, training the routing network (or router) in MoE architectures introduces the challenge of optimizing a non-differentiable, discrete objective (Shazeer et al., 2017; Zoph et al., 2022). Various techniques, such as switch routing (Fedus et al., 2022), top-$k$ expert-choice routing (Zhou et al., 2022), and linear programming (Lewis et al., 2021), have been developed to address this challenge, often requiring carefully designed load-balancing objectives (Fedus et al., 2022) or introducing additional complexity in assignment algorithms (Lewis et al., 2021; Roller et al., 2021).
Recent research is exploring fully differentiable MoE architectures to overcome training challenges. SMEAR (Muqeeth et al., 2023), for instance, softly merges experts by averaging their parameters rather than activating the top-$k$ experts. However, SMEAR’s effectiveness has only been shown in small-scale fine-tuning experiments on downstream classification tasks (Wang et al., 2018).
@@
In this work, we propose Lory [^1] Lory is a tribe of parrots known for their rainbow-like colors, reflecting the spirit of ’soft’ MoE., the first approach that scales such fully differentiable MoE architectures to autoregressive language model pre-training. Unlike text classification tasks, which only require routing each input sequence to different experts, language modeling makes predictions for each input token. Performing token-level routing is prohibitively expensive because the computational cost of merging operations scales linearly with the number of experts.
@@
Lory is based on two key techniques (Figure 1). First, we propose causal segment routing: for a sequence of input tokens, we split them into multiple segments of fixed length and use the previous segment to determine the router’s weights and calculate the merged expert for the subsequent segment. During inference, we can simply use the prompt to make a single routing decision throughout the generation process. This segment-level routing strategy preserves the autoregressive nature of language models while keeping the merging operations efficient. However, since the text data for pre-training language models usually concatenates random sets of documents, we find that such routing can lead to scenarios in which experts are not sufficiently specialized. To address this, we propose our second technique: similarity-based data batching for MoE training, which groups semantically similar documents to form consecutive segments. This idea has been recently proposed to train LMs to better reason across document boundaries (Shi et al., 2024), and we find that it leads to more effective training of expert routing.
We pre-train a series of Lory models from scratch under a training budget of 150B tokens, with 0.3B and 1.5B active parameters, and 8, 16, or 32 experts (corresponding to 6.8B and 29.5B full parameters at most; see Table 4). Experimental results show that our Lory models significantly outperform equal-sized dense models trained with the same amount of data, achieving performance gains in both perplexity ($+13.9\%$) and a wide range of downstream tasks, including commonsense reasoning ($+3.7\%$), reading comprehension ($+3.3\%$), closed-book QA ($+1.5\%$), and text classification ($+11.1\%$). Interestingly, despite using segment-level routing, Lory achieves competitive performance compared to state-of-the-art MoE models with token-level, non-differentiable discrete routing (Zhou et al., 2022). Our analysis further shows that the trained experts capture domain-level specialization without any supervision, making it distinct from previous MoE LMs with token-level routing, which only exhibit local patterns uniformly distributed across different domains (Xue et al., 2024; Jiang et al., 2024). Together, we present the first fully differentiable MoE model suitable for language model pre-training and demonstrate its effectiveness at scale. We hope our work sheds light on the potential of fully differentiable MoE architectures in cultivating specialized experts and encourages continued exploration in this research field.
@@ 2.1 Sparsely-activated MoE
Transformer-based MoE language models typically substitute feed-forward network (FFN) layers with sparsely activated MoE layers (Shazeer et al., 2017; Fedus et al., 2022; Zoph et al., 2022). Assume an MoE layer consists of $E$ expert FFNs, each parameterized as $\text{FFN}(\cdot;\theta_{1}),\dots,\text{FFN}(\cdot;\theta_{E})$, where the function $\text{FFN}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ defines a single expert module. For each token $x$ in a sequence, an MoE layer takes the hidden representation $h_{x}\in\mathbb{R}^{d}$ as the input and computes its output $o_{x}\in\mathbb{R}^{d}$ by sparsely activating $k$ experts in this layer and aggregating the outputs through a weighted sum:
@@
The routing weight $e_{i}$ for the $i$-th expert is determined by a routing network or router $R$, which takes $h_{x}$ as input and calculates the weight for each expert. In practice, to achieve sparsity and computational efficiency, only one (Fedus et al., 2022) or $k$ (Lepikhin et al., 2021) experts with the highest routing weights are activated at each MoE layer. The weights of the remaining experts are set to 0 (i.e., $e_{i}=0$), eliminating the need to compute $\text{FFN}(h_{x};\theta_{i})$ and effectively deactivating the $i$-th expert.
@@ 2.2 Fully Differentiable MoE Architectures via Expert Merging
The primary challenges in training sparsely activated MoE models arise from the difficulty in training discrete routers. A promising direction is to design fully differentiable MoE architectures that do not rely on extra loss formulations for stabilized training. A recent model architecture (Muqeeth et al., 2023) demonstrates the feasibility of this approach by computing a weighted average of all expert FFNs in the parameter space (Matena & Raffel, 2022; Wortsman et al., 2022), thereby creating a “merged FFN.” Given an input $x$ and its corresponding routing weights $e_{i}$, the output $o_{x}$ of a merged FFN is computed as:
@@
However, naively extending this approach to autoregressive language models, which would require computing the merged FFN for each token in a sequence, is infeasible as the computational costs of merging operations scale linearly with the number of experts. SMEAR (Muqeeth et al., 2023) has only been evaluated for downstream fine-tuning on text classification tasks, making routing decisions based on a pooling representation of the entire input sequence, i.e., $e_{i}=\text{Softmax}(R(\frac{\sum_{j=1}^{L}h_{x_{j}}}{L}))_{i}$. Such operations disrupt the autoregressive property in language model pre-training. In this work, we address these challenges by developing a fully differentiable MoE architecture suitable for autoregressive language modeling and pre-training such models at scale.
@@
Lory is an approach for pre-training fully differentiable MoE language models (Figure 1). The core technique that enables Lory to be fully differentiable is expert merging (Muqeeth et al., 2023, see details in Section 2.2). To make it computationally feasible, we propose a causal segment routing method that merges experts only once per segment, effectively reducing the number of merging operations (Section 3.1). We also propose a data batching strategy that groups semantically similar texts, which is crucial for the effective training of the segment-level router (Section 3.2).
@@
We denote an input sequence of $L$ tokens as $X=(x_{1},x_{2},\ldots,x_{L})$. Considering a segment size $T$, we divide the input sequence into $N=\lceil L/T\rceil$ segments, denoted as $S_{1},S_{2},\ldots,S_{N}$. We use $R$ to denote the routing network (parameterized as a linear layer), which computes the weights for expert merging. Let $h_{x}$ represent the hidden representation of the token $x$. The parameters of the $i$-th expert FFN are denoted by $\theta_{i}$.
@@ 3.1 Efficient Expert Merging via Causal Segment Routing
An intuitive way to reduce the computational cost is to use segment-level routing instead of token-level routing, which can reduce the number of merging operations from $L$ to $N$. However, simply using the current segment to compute the routing weights can cause information leakage.
@@
We propose causal segment routing to effectively route information across segments in an autoregressive manner. [^2] Pseudocode of the causal segment routing strategy can be found in Appendix A. This method merges FFNs in an MoE layer based on the previous segment’s information and uses it to process the current segment. Specifically, given a training instance $X$ that consists of $L$ tokens (e.g., $L=4096$), we split the training instance into $N$ segments, each containing $T$ (e.g., $T=256$) consecutive tokens. For the $k$-th segment $S_{k}$, where $k>1$, we compute the average of the hidden representations of its preceding segment $S_{k-1}$, denoted as $\bar{h}_{k-1}$. Using the average hidden representation allows the model to adapt to prompts of varying lengths during inference. $\bar{h}_{k-1}$ is then utilized to determine the routing weights, resulting in a merged expert $\bar{\theta}$:
@@
We then use the merged expert $\bar{\theta}$ to process all the tokens in the current segment $S_{k}$, i.e., $o_{x}=\text{FFN}(h_{x};\bar{\theta}),\forall x\in S_{k}$. This approach ensures that the routing decisions made by the model are based exclusively on data from preceding positions. For the first segment $S_{1}$, the representation of the segment itself is used to compute the merging weights for its own FFN. To prevent information leakage, we implement a stop-gradient operation on $R(\bar{h}_{1})$. As demonstrated in Appendix B, merging experts at the segment level incurs minimal overhead compared to the training of dense models.
@@
During inference, we begin with a given prompt and make a single routing decision per layer based on the average hidden representations of the prompt. This routing decision determines a merged FFN, which is used consistently throughout the entire generation process. It is important to note that this inference process is as simple and computationally efficient as that of dense models. [^3] In Appendix G.2, we compare the prompt-only routing strategy to using the causal segment routing strategy that faithfully follows the training design, and find they do not lead to significant differences. We also discuss the potential of converting Lory to sparsely MoE models for memory-efficient inference.
@@ 3.2 Similarity-based Data Batching
The standard practice of pre-training LMs is to randomly concatenate documents to construct training instances with a fixed length. This approach can lead to under-specialized experts because tokens within adjacent segments may come from very different and irrelevant documents.
To mitigate this issue, we employ a similarity-based data batching technique inspired by Shi et al. (2024), which sequentially concatenates similar documents to construct training instances. This method encourages high similarity between adjacent segments, enabling the experts to specialize in different domains or topics.
We use Contriever (Izacard et al., 2022) to measure document similarity and apply a greedy search algorithm to concatenate similar documents to form batches (see Appendix C). Although our data batching technique is similar to that of Shi et al. (2024), our goal is different. While they focus on enhancing language models’ reasoning across document boundaries, we find this approach particularly effective for fostering expert specialization during MoE model training.
@@ 4.1 Setups
We employ the AdamW optimizer (Loshchilov & Hutter, 2019) with $\beta_{1}=0.9$ and $\beta_{2}=0.95$ and use a learning rate of 2e-4 with a cosine learning rate scheduler. All models with a batch size of 1 million tokens. We employ the data parallelism with the ZeRO optimization (Rajbhandari et al., 2020) for distributed training. At the beginning of training, we train a parameter-matched dense model and duplicate the FFN layers as initialization of the MoE model. In our experiments, we use the first 5% training steps as the warmup to initialize the MoE weights. We find that without warmup training, there may be more experts under-utilized (see Appendix G.3 for an ablation study). We also apply a linear warmup to the learning rate scheduler for the first $5\%$ training steps. We train our models with up to 64 A100 GPUs.
@@
We randomly sample a subset (150 billion tokens) of the Commoncrawl dataset (Wenzek et al., 2019) for training. Using the similarity-based data batching method from Shi et al. (2024), we construct all training instances (see Appendix C for details).